# HiCL: Hippocampal-Inspired Continual Learning

### v1

@@ Discussion
This approach drastically lowers the total computational footprint. For instance, on Split CIFAR-10, our Small model requires 10.81 MFLOPs for the shared backbone and 5.48 MFLOPs for a single expert, resulting in an efficient inference cost of only 16.3 MFLOPs. The Large model’s cost under the same conditions is 49.7 MFLOPs. On the more complex Split Tiny-ImageNet, these costs are 49.1 MFLOPs for the Small model and 172.7 MFLOPs for the Large model. While activating a single expert may offer less GPU parallelism than a batch-parallel approach, the significant reduction in total operations makes HiCL well-suited for continual learning on resource-constrained hardware.
Mechanistic Insights
The core of HiCL’s success lies in its DG‑based gating mechanism. By enforcing top‑$k$ sparsity (Equation 1) and aggregating DG outputs into per‐expert prototypes via an EMA update, each input is mapped to a near‐orthogonal code subspace [7]. At inference, cosine similarity between live codes and prototypes yields routing accuracy exceeding 95%, consistent with high‐dimensional geometry results for random sparse vectors [54]. This sparse separation ensures expert misrouting is rare and each expert specializes effectively on its assigned tasks.
@@
The success of our routing mechanism stems from the fact that each expert’s Dentate Gyrus (DG) module learns a unique “feature vocabulary” tailored to its specific task. For an expert trained on Task 1 (e.g., distinguishing cats from dogs), its DG module becomes highly attuned to features like fur texture, whiskers, and ear shape. In contrast, an expert trained on Task 2 (e.g., airplanes vs. ships) will learn to amplify features related to wings, fuselages, and water.
@@
This separation is quantified by the prototype similarity matrices in Figure 4b and 4d. The strong diagonal shows that each prototype is similar to itself, while the low off-diagonal values (all close to zero) confirm that the prototypes for different tasks are nearly orthogonal. Therefore, the routing decision, a simple cosine similarity check against the prototypes is not just a superficial pattern match. It is a measure of representational resonance. The input is naturally routed to the expert whose internal feature vocabulary is best suited to understand and process it, a choice made simple and reliable by the clear, quantifiable separation of the learned feature spaces.
@@
Despite its strengths, HiCL has several limitations. First, it operates under a Task IL assumption, requiring known task boundaries and labels during training. While Class IL inference is label free, Phase 1 still depends on task identities; fully unsupervised gating remains an open direction. Second, we use $\rho=0.05$ and $N=5/10$ experts across all tasks. . These hyperparameters may not generalize well to domains with different levels of task heterogeneity or complexity. Third, routing can be sensitive in rare cases, with less than five percent of samples showing class confusion due to overlapping DG codes. This could potentially be addressed with adaptive gating thresholds or multi prototype routing. Lastly, our experiments are limited to vision datasets with LeNet style backbones. The model’s performance on larger scale datasets or nonvisual domains such as language or reinforcement learning remains unexplored.
@@
Broader Impact
HiCL’s demonstration that sparse hippocampal motifs can yield efficient, robust continual learners opens pathways for real‐world AI systems that must learn adaptively under resource constraints—such as autonomous robots, edge computing devices, and lifelong learning agents. By bridging neuroscience and machine learning, we hope to inspire architectures that are not only performant but also interpretable, modular, and biologically grounded.

### v2

@@ Discussion
This approach drastically lowers the total computational footprint. For instance, on Split CIFAR-10, our Small model requires 10.81 MFLOPs for the shared backbone and 5.48 MFLOPs for a single expert, resulting in an efficient inference cost of only 16.3 MFLOPs. The Large model’s cost under the same conditions is 49.7 MFLOPs. On the more complex Split Tiny-ImageNet, these costs are 49.1 MFLOPs for the Small model and 172.7 MFLOPs for the Large model. While activating a single expert may offer less GPU parallelism than a batch-parallel approach, the significant reduction in total operations makes HiCL well-suited for continual learning on resource-constrained hardware. Table 1 reports the full model capacity for fair comparison with baselines that execute all parameters, while conditional computation enables the 16.3 MFLOPs efficiency advantage described above.
@@
The success of our routing mechanism stems from effective pattern separation in the DG layer, as evidenced by our ablation study (Table 2). Removing DG gating or the contrastive loss collapses Class-IL accuracy from 57.99 to approximately 19, while omitting top-$k$ sparsity reduces Task-IL accuracy from 92.67 to 55.85, demonstrating that sparse separation is critical to the model’s performance.
Each expert’s Dentate Gyrus (DG) module learns a unique feature vocabulary tailored to its specific task. For an expert trained on Task 1 (e.g., distinguishing cats from dogs), its DG module becomes highly attuned to features like fur texture, whiskers, and ear shape. In contrast, an expert trained on Task 2 (e.g., airplanes vs. ships) will learn to amplify features related to wings, fuselages, and water.
@@
This separation is quantified by the prototype similarity matrices in Figure 4b and 4d. The strong diagonal shows that each prototype is similar to itself, while the low off-diagonal values (all close to zero) confirm that the prototypes for different tasks are nearly orthogonal. Therefore, the routing decision, a simple cosine similarity check against the prototypes is not just a superficial pattern match. It is a measure of representational resonance. The input is naturally routed to the expert whose internal feature vocabulary is best suited to understand and process it, a choice made simple and reliable by the clear, quantifiable separation.
@@
Configuration Task-IL Class-IL
Full HiCL 92.67 57.99
   w/o Contrastive 91.56 19.19
   w/o DG Gating 91.82 18.39
   w/o top k Sparsity 55.85 30.71
   w/o CA1 92.64 52.39
   w/o EMA 90.63 56.29
   w/o CA3 92.29 58.73
   w/o EWC 89.48 58.92
Table 2: Ablation study on Split CIFAR-10 (Replay szie = 200)
@@
Despite its strengths, HiCL has several limitations. First, it operates under a Task IL assumption, requiring known task boundaries and labels during training. While Class IL inference is label free, Phase 1 still depends on task identities; fully unsupervised gating remains an open direction. Second, we use $\rho=0.05$ and $N=5/10$ experts across all tasks. These hyperparameters may not generalize well to domains with different levels of task heterogeneity or complexity. Third, routing can be sensitive in rare cases, with less than five percent of samples showing class confusion due to overlapping DG codes. This could potentially be addressed with adaptive gating thresholds or multi prototype routing. Lastly, our experiments are limited to vision datasets with LeNet style backbones. The model’s performance on larger scale datasets or nonvisual domains such as language or reinforcement learning remains unexplored.